OpenStudy (anonymous):

a rectangles lenght is 4 inches more than its width. the sum of the length and width is 13 inches. what is the width?

OpenStudy (anonymous):

width is nine.

OpenStudy (anonymous):

How do you get this?

OpenStudy (anonymous):

well, the sum means they are added together, so L+W=13. And you know that L=w+4. So plug back is L to get (w+4) + w=13, subtract four from each isde 2w=9, so my bad! w=4.5

OpenStudy (anonymous):

want to help with a few more?

OpenStudy (anonymous):

sure, does that explanation make sense though?

OpenStudy (anonymous):

yes.. I am trying to help my son with his homework... I never had algebra.. Thanks!!

OpenStudy (anonymous):

haha :) sons dont know math ;)

OpenStudy (anonymous):

mike worked 2 more hours on sat. than on fri. He worked 15 hours in the two days. How long did he work on Sat?

OpenStudy (anonymous):

He actually is smart, but the teacher just handed this to the kids on the way out of class..I really appreciate this!

OpenStudy (anonymous):

it would be set up the same way. so the equation would look like: (x+2) +x=15 x being satuday. so then 2x=13 divide by 2 and get 6.5 :)

OpenStudy (anonymous):

haha, yeah i had my fair share of those papers :)

OpenStudy (anonymous):

How old are you?

OpenStudy (anonymous):

i am 15 :)

OpenStudy (anonymous):

OMG smart girl!!!

OpenStudy (anonymous):

thank you :) at least with math that is. i had algebra 1 in 7th grade(i am currently in 10th)

OpenStudy (anonymous):

Ann and Tom saw two movies in one day. The two movies lasted 4 hours. one movie was half an hour long then the other. How many hours did the shorter movie run?

OpenStudy (anonymous):

Where do you go to school? Did you test out to be in a higher grade?

OpenStudy (anonymous):

same as the last ones :) you sent it up as a formula. M+M+1/2=4

OpenStudy (anonymous):

I started online schooling this year, and no i just took higher classes.

OpenStudy (anonymous):

So is it safe to say they will all be set up the same? Awesome for you!!!

OpenStudy (anonymous):

if they are all written as that one is longer than the other or one is the other plus this then yes! :D

OpenStudy (anonymous):

Your funny! What are you going to be when you grow up? He came up with an answer of 1 3/4.

OpenStudy (anonymous):

haha, why thank you :) i try! and i have no clue, ive been thinking mechanical engineer because that kind of stuff really interests me. and that is right! tell him good job :)

OpenStudy (anonymous):

Very cool!! he is stumped on this one tho...just give us the set up again...

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

the higher temperature one day was twice the low temp. the high temp was also 33 degrees more than the low temp. what was the low temp?

OpenStudy (anonymous):

this would be a multiplication one, not addition like the others.

OpenStudy (anonymous):

hmm.. Im not gettin that... and Im old!! and in college!!

OpenStudy (anonymous):

haha its fine :) it is saying the high temp in twice the low temp which means H=2L the high temp was also 33 more than the low temp H=L+33 So setting them equal to each other L+33=2L then i would subtract L to get 33=L So then plug back in L h=(33)+33 so the high temp would be 66

OpenStudy (anonymous):

I dont need math for counseling!! Thankfully! Thank you!

OpenStudy (anonymous):

haha, thats always good :) thats how i feel about history!

OpenStudy (anonymous):

if you ever need someone to help just message me or mention me and i will try my best to help out!

OpenStudy (anonymous):

Some hikers made the round trip on a mountain trail in 10 hours. It took three times longer to hike up than to hike down. How many hours did it take to hike up the trail?

OpenStudy (anonymous):

same as the one before, it would be a multiplication one. and plugging it back in

OpenStudy (anonymous):

Maybe I should have my son just add this page and your name? He is 14, and gets all A's but was very confused... ok.. Thanks again for your help!!! Is there a way to save you?

OpenStudy (anonymous):

there are security settings you can put on his as well, such as people no tbeing able to message him if he is not a "fan", and if you click on my name and at the bottom where it says become a fan it will save me :)

OpenStudy (anonymous):

awesome thanks!!

OpenStudy (anonymous):

yupp :)

OpenStudy (anonymous):

I am now your fan! Thanks again and keep up the great work!!!

OpenStudy (anonymous):

yupp! and i will do so :) thank you

OpenStudy (anonymous):

:o)

OpenStudy (anonymous):

are you still here?

OpenStudy (anonymous):

i am now, need more help?